OpenStudy (anonymous):

Limits Calculus 1) find slope of the tangent line to the graph of f at the given point. 2) find the slope-intercept equation of the tangent line to the graph of f at the given point. Problem #1: f(x)=-3x+4 at (1,1) Problem #2: f(x)=3x^2 at (1,3) Please help me, I found the answers but I'm not sure if they are right :\ THANKS!

OpenStudy (anonymous):

for #1 \(f(x)=3x+4\) you have a line the slope is a constant, namely \(3\)

OpenStudy (anonymous):

doesn't matter what the point is for the first one, the slope is always 3

OpenStudy (anonymous):

for the second one if \[f(x)=3x^2\] then \[f'(x)=6x\] if you want the slope at \(x=1\) find \[f'(1)\]

OpenStudy (anonymous):

I know but I don't think I got the right answer.

OpenStudy (anonymous):

at a point of figuring, I got -3-3h+4-1

OpenStudy (anonymous):

for the first one or the second one?

OpenStudy (anonymous):

First

OpenStudy (anonymous):

the first one is a line you do not need a limit the slope of a line is constant

OpenStudy (anonymous):

if you are trying to compute \[\frac{f(1+h)-f(1)}{h}\] we can do that too

OpenStudy (anonymous):

oh and you get exactly what you wrote only divided by \(h\)

OpenStudy (anonymous):

\[\frac{ -3-3h+4-1}{h}=\frac{-3h}{h}=-3\]

OpenStudy (anonymous):

Oh I see!

OpenStudy (anonymous):

but in any case it is the slope of \(y=-3x+4\) which is \(-3\)

OpenStudy (anonymous):

For the second one, I did 3(1+h)^2-3(1)^2/h and I got 3(1+2h+h^2)-9 then 3+6h+3h2-9 is that right?

OpenStudy (anonymous):

3h^2 I mean

OpenStudy (anonymous):

\(3\times 1^2=3\) not \(9\)

OpenStudy (anonymous):

numerator should be \[ 3+6h+3h2-3\]

OpenStudy (anonymous):

which is \[6h+3h^2\]

OpenStudy (anonymous):

divide by \(h\) and you get \[6+3h\] then let \(h\to 0\) and get \(6\)

OpenStudy (anonymous):

Ah, I messed up on the 3(1)^2 and didn't notice my mistake. Thank you so much for helping! Medal for you!

OpenStudy (anonymous):

yw